# Misinformation is not about Bad Facts: An Analysis of the Production and Consumption of Fringe Content

### v1

@@ Results.
Table 3 presents the 21 pairs (LIWC category, article topic) for which the difference between the far-right publishers and the moderates is statistically significant – 14 LIWC categories for ‘Top stories,’ 5 for ‘Australia’ and one for each of ‘Finance’ and ‘Climate change.’ However, most of these have low ($d\sim 0.2$) or moderate ($d\sim 0.5$) effect sizes; only two such topic-category pairs show large effect sizes ($d\sim 0.8$). In ‘Finance’, far-right publishers use significantly more words from the male category – containing 230 words such as ‘he,’ ‘his,’ ‘him’ or ‘man’ ($d=1.01$). The second is ‘Climate change’, with far-right articles using statistically significantly more words in the Negative Emotions category (618 words such as ‘bad,’ ‘hate,’ ‘hurt,’ ‘worry,’ ‘fear’) and the effect size is large ($d=1.03$).
@@ Conclusion.
The linguistic signals captured by LIWC alone are insufficient for distinguishing extreme (far-right) publishers from moderate ones. Corroborated with the conclusion from Section 3.1, this indicates little difference in what publishers of different political leanings produce.
3.3 Far-right articles: production vs. consumption
@@
Here, we assess whether there is a difference in the linguistic styles of the articles far-right users share. Note there is a difference with the analysis in the previous Section 3.2: there we analyzed the difference in information production by far-right publishers, here we investigate information consumption and sharing by far-right users.
@@
The far-right Twitter users share $11,643$ articles from the Google News dataset on the four topics of interest – ‘Top Stories,’ ‘Australia,’ ‘Finance,’ and ‘Climate Change’. We process these articles’ content using LIWC. Finally, we perform t-tests (Bonferroni corrected) to compare articles shared by the far-right Twitter users with 1) articles produced by far-right publishers and 2) articles produced by moderate publishers. Table 4 shows select LIWC categories that show statistically significant differences.
@@ Results.
There are several noteworthy findings. First, we find $64$ (LIWC category, article topic) pairs statistically significantly different when comparing articles from far-right producers with articles shared by far-right Twitter users. In comparison, there are $325$ significantly different pairs between articles from moderate publishers and those shared by far-right Twitter users. However, only $21$ pairs significantly differ between far-right and moderate publishers (see Section 3.2). This indicates that far-right publishers are more similar to moderate publishers than to those far-right Twitter users share. Investigating the mean values ($\mu$) and standard deviation ($\sigma$) yields a similar conclusion – the differences between what the far-right Twitter users chose to share and what is produced by the far-right publisers are larger the differences between far-right and moderate publishers. Second, the LIWC categories that showed significant differences were Words Per Sentence (WPS), Culture, and Politic, where the articles shared by the Twitter users showed significantly higher mean values. ‘Climate change’ did not show significant results in any LIWC category. We also report Lifestyle and money, which showed significant differences in ‘Top stories’ and ‘Australia’ with high effect sizes. Third, the articles shared by Twitter users use more words from the categories of Culture and, specifically, its subcategory, politic, compared to the far-right-leaning articles from the Google news dataset. This indicates that far-right Twitter users share more political articles than the general Google News sample suggests. This seems intuitive, as most conspiracy theories touch on politics.
@@ Conclusion.
The above results suggest that information consumption patterns of online far-right users differ from the information production by far-right publishers. We hypothesize that these users do not share random samples of the articles produced by far-right publishers. They rather cherrypick the articles most useful for their arguments. We note that users may not necessarily express agreement or a positive point of view of the article that they share. In fact, far-right Twitter users regularly share articles from reputable and left-leaning publishers. Many of these articles are about politics, so negative responses may be anticipated.
4 Style over Content
In this section, we propose a style classifier and attempt to distinguish different online extreme groups (Section 4.1), styles (Section 4.2), and producing from sharing (Section 4.3).
4.1 Identify Extreme Groups using Styles (RQ2)
Lee et al. (2022) showed that political groups with left- and right-leaning ideologies exhibit distinct tendencies when consuming and disseminating information on social media platforms. We curated two collections of Facebook groups and collected their activity from over two years, presenting two datasets of extreme groups – antivax and far-right, as detailed in Section 2.
Why style?
Style words reflect how people communicate, whereas content words convey what they say in terms of meaning and topic. Style words are more closely linked to measures of people’s social and psychological worlds (Tausczik and Pennebaker 2010). Styles encompass a range of linguistic features, including sentence structure, grammar, and punctuation patterns. Unlike the content, which can be influenced by subject matter, external sources, or intentional deception, stylistic features are intrinsic to one’s writing and are less prone to deliberate manipulation.
We focus on stylistic differences because a group’s contents or topics of interest can dynamically change; however, stylistic fingerprints persist (da Rocha et al. 2022). Another reason to consider style over content-related features is that stylistic representation is applicable across domains, while content-related features are constrained to specific topics.
@@ 4.1 Identify Extreme Groups using Styles (RQ2)
We begin by demonstrating the classification performance in distinguishing between extreme groups. Subsequently, we differentiate various text styles, irrespective of their association with specific groups.
We design a predictive experiment aimed at evaluating and contrasting the effectiveness of three dictionary-based stylistic metrics (LIWC, Grievance and StyloMetrix). The task is to distinguish posts from the two extreme Facebook groups and a normal (non-extreme) group – we created the normal group to encompass a wide range of discussions, from cooking to non-profit organizations. To put the performance of stylistic classifiers into context, we compare against a content-based baseline – the popular text encoding technique, BERT (Devlin et al. 2018) commonly used in textual classification tasks.
In contrast to the three dictionary-based encodings, BERT also considers the contextual information for each instance of a given word, enhancing its capabilities. The BERT embeddings used in this study are the mean-pooled representations of the activations from the last layer of the en_stsb_distilbert_base model. This approach captures contextual information from the entire input sentence and produces fixed-length vectors suitable for downstream tasks.
Table 5: Performance of five feature sets. We distinguish between two extreme groups – Far right and Antivax – and the normal Facebook group. We randomly sample $1,000$ posts for each group. We report accuracy and macro-F1 to compare among the classifiers and features used.
@@ Results.
Since BERT leverages the content and style of the given text, it is often the best-performing feature set. However, stylistic-based classifiers outperform the content-based classifiers: using the Random Forest classifier, LGS outperforms BERT. In other words, based solely on the style and without leveraging the content of a given post, we can predict which ideological group the post came from. The true positive rate for the normal group was $0.95$, for Far right group is $0.79$, and for Antivax group is $0.66$. We notice that some posts from Antivax group were misclassified as Far right, while only few posts from Far right were misclassified as Antivax. This is because anti-vax space is a unique niche with far-right elements, while far-right groups generally discuss diverse topics.
@@ 4.2 Identifying Styles
Here, we classify the writing styles of posts. This predictive exercise differs from the one in Section 4.1; here, we test whether human-labeled styles can be detected using a style-based classifier; we distinguished user groups in Section 4.1.
@@
First, a research team member with expertise in literature science manually annotated the writing style of $100$ text samples from the Far right and Antivax groups ($50$ for each group). They employed a deep qualitative approach in which the style labels were generated alongside the labeling. We identified styles that are extensively used in misinformation online environments by manually inspecting them. The labeled styles in these samples are “Casual”, “Empowerment”, “Clickbait”, “Expert” and “Intimacy”. This annotation is a result of sampling 50 posts from both Far right and Antivax then We remove the posts with no style; Table 6 shows the number of posts per style.
@@
Table 7: Classification results of the three styles; “Casual”, “Empowerment” and “Expert”. A Random Forest classifier and a Logistic Regression was used with a stratified split. The results show the average performance of 2-fold cross-validation due to the size of the samples.
@@
Empowerment Random Forest 0.60 0.58 0.51 0.54
Logistic Regression 0.49 0.52 0.67 0.66
Casual Random Forest 0.67 0.5 0.54 0.55
Logistic Regression 0.45 0.52 0.67 0.70
Expert Random Forest 0.47 0.5 0.42 0.5
Logistic Regression 0.40 0.73 0.57 0.46
“Clickbait” and “Intimacy” have only two and one exemplars, respectively, so we remove them and classify “Casual”, “Empowerment” and “Expert” styles. For each of the style, we performed a binary classification by randomly sampling negative samples to balance the sample size. For the Random Forest classifier, we used max tree depth = $3$ and the number of trees = $8$ due to the small sample size.
@@ Results.
Table 7 reports the classification results. In general, the stylistic features, LGS, perform better with the Random Forest classifier and BERT shows better performance with the Logistic Regression which is expected due to the nature of these features. While it is true that the Random Forest classifier may not fully leverage the rich contextual information in BERT embeddings, we aimed to assess whether it could still yield valuable insights or competitive results when compared to more complex models. This experimental choice was made to offer a comparative perspective on different feature representations and classifiers.
For all three styles, the LGS stylistic features outperform BERT with the Random Forest classifier measured by macro F1 score, which is consistent with the results presented in Table 5 for group style classification. “Expert” style was the most challenging for both classifiers since there were only $13$ samples.
4.3 Identify Far-right articles using Styles: production vs. consumption (RQ3)
Inspired by the findings in Section 3.3, we hypothesise that far-right production can be differentiated from far-right consumption. In this section, we the style classifier to distinguish far-right production from far-right consumption. We add moderate production for comparison purposes.
@@
We use $155,669$ Twitter shared articles explained in Section 2 as far-right consumption data since these were shared by far-right Twitter users. In order to increase the number of articles produced by far-right publishers, as well as moderate, we expand the number of articles as follows: for each article shared by Twitter users if the publisher of the article is referenced in allsides data, we add this article to far-right production or moderate production based on the leaning of the publisher. After the expansion, we finally have $155,669$ Twitter shared articles (consumption), $41,995$ moderate articles and $14,758$ far-right produced articles. To balance the sample size among the three classes, we randomly sample $14,758$ articles from Twitter shared and moderate produced articles so that all three classes have $14,758$ samples.
@@
The reported scores in Table 8 are averaged over 10-fold stratified cross-validation.
Table 8: Classification results of the far-right produced articles, moderate articles, and the articles consumed by far-right users. A Random Forest classifier and a Logistic Regression was used to report the performance of 10-fold stratified cross-validation.
Classifier LGS BERT
macro F1 Accuracy macro F1 Accuracy
consumption Random Forest 0.65 0.65 0.40 0.54
Logistic Regression 0.56 0.59 0.50 0.60
FR production Random Forest 0.75 0.65 0.64 0.54
Logistic Regression 0.70 0.59 0.71 0.60
MD production Random Forest 0.53 0.65 0.52 0.54
Logistic Regression 0.48 0.59 0.58 0.60
In Section 4.2, we observed that our curated stylistic features (LGS) perform better with Random Forest classifier while BERT could outperform with Logistic Regression in most cases. In Table 8, we see that LGS with Random Forest classifier consistently outperforms BERT with Logistic Regression except for articles from moderate publishers. In addition, even when LGS with Random Forest classifier did not outperform BERT with Logistic Regression, the performance difference was minimal as witnessed by the case of far-right produced articles.
Figure 2: Confusion matrix of the three classes from Random Forest classifier.
Fig. 2 reports the confusion matrix of the three class classification from Random Forest classifier. This result is the average of 10-fold stratified cross-validation. The most notable observation is that far-right produced articles are easily distinguished from the others as well as far-right consumption. The most number of confusions occurred between the articles between the far-right consumption and the moderate produced. These findings reinforce the previous observations from Section 3.3 that far-right production and far-right consumption patterns are different.

### v2

@@ Results.
Table 3 presents the 21 pairs (LIWC category, article topic) for which the difference between the far-right publishers and the moderates is statistically significant – 14 LIWC categories for ‘Top stories,’ 5 for ‘Australia’ and one for each of ‘Finance’ and ‘Climate change.’ However, most of these have low ($d\sim 0.2$) or moderate ($d\sim 0.5$) effect sizes; only two such topic-category pairs show large effect sizes ($d\gtrsim 0.8$). In ‘Finance’, far-right publishers use significantly more words from the male category – containing $230$ words such as ‘he,’ ‘his,’ ‘him’ or ‘man’ ($d=1.01$). The second is ‘Climate change’, with far-right articles using statistically significantly more words in the Negative Emotions category – $618$ words such as ‘bad,’ ‘hate,’ ‘hurt,’ ‘worry,’ ‘fear’ ($d=1.03$).
@@ Conclusion.
The linguistic signals captured by LIWC alone are insufficient for distinguishing extreme (far-right) publishers from moderate ones. Corroborated with the conclusion from Section 3.1, this indicates little difference in the content and topic of what publishers of different political leanings produce overall.
3.3 Far-right articles: content production vs consumption
@@
Here, we assess whether there is a difference in linguistic patterns based on content in the articles that far-right users share. Note there is a difference with the analysis in the previous Section 3.2 in which we analyzed the difference in information production by far-right publishers, here we investigate information consumption and sharing by far-right users.
@@
The far-right Twitter users shared $11,643$ articles from the Google News dataset on the four topics of interest – ‘Top Stories,’ ‘Australia,’ ‘Finance,’ and ‘Climate Change’. We process these articles’ content using LIWC. Finally, we perform t-tests (Bonferroni corrected) to compare articles shared by the far-right Twitter users with 1) articles produced by far-right publishers and 2) articles produced by moderate publishers. Table 4 shows select LIWC categories that show statistically significant differences. There are several noteworthy findings which we introduce below.
@@ Results.
First, we find $64$ (LIWC category, article topic) pairs statistically significantly different when comparing articles from far-right producers with articles shared by far-right Twitter users. In comparison, there are $325$ significantly different pairs between articles from moderate publishers and those shared by far-right Twitter users. However, only $21$ pairs significantly differ between far-right and moderate publishers (see Section 3.2). This indicates that the overall output of far-right publishers is generally more similar to moderate publishers in content-based linguistic patterns, however the far-right Twitter users share articles that exhibit different and less common linguistic patterns.
Investigating the mean values ($\mu$) and standard deviation ($\sigma$) yields a similar conclusion – the differences between what the far-right Twitter users chose to share and what is produced by the far-right publishers are larger than the differences between far-right and moderate publishers. Second, the LIWC categories for the articles that the far-right users shared showed significant differences in the categories Words Per Sentence (WPS), Culture, and politic, with the articles Twitter shared showing significantly higher mean values. The Culture category includes words relating to nations, political processes (politic), and ethnic identities. Lifestyle and money also showed significant differences in the topics of ‘Top stories’ and ‘Australia’ with high effect sizes. Lifestyle includes words that discuss money, households, employment, and religion. This indicates that far-right users typically shared articles that were about politics, societal make-up, and money – common key issues for right-aligned groups. Third, the articles shared by Twitter users use more words from the categories of Culture and, specifically, its subcategory, politic, compared to the far-right-leaning articles from the Google news dataset. This indicates that far-right Twitter users share more political articles than the general Google News sample suggests. This seems intuitive, as most conspiracy theories touch on politics.
@@ Conclusion.
The above results suggest that the information consumption patterns of online far-right users differ from the scope of article production by far-right publishers. We hypothesize that these users do not share random samples of the articles produced by far-right publishers. Rather, they rather cherry pick the articles most useful for their arguments. We note that users may not necessarily express agreement or a positive point of view of the article that they share. In fact, far-right Twitter users regularly share articles from reputable and left-leaning publishers, potentially as evidence for a circumstance they wish to condemn. This suggests that far-right users are willing to draw on any news articles that evidence their viewpoints, regardless of the political leaning of the source (although they seem to prefer far-right over moderate produced articles, see Figure 2). This is quite different from what has been proposed by previous research, which is that users’ engagement with a news article is heavily mediated by the source it is attributed to, and whether the user has an overall positive or negative view of the news media company itself (Yun et al. 2018).
4 The Style: Style over Content
In the previous sections, we learned that the integrity of contents do not have strong association with the ideologies of the publishers. Additionally, contents produced by far-right publishers and consumed by far-right users are different which was unexpected. In this section, instead of comparing contents, we investigate the style of texts from different groups and distinguish between them.
Why style over content?
Style words reflect how people communicate, whereas content words convey what they say in terms of meaning and topic. Style words are more closely linked to measures of people’s social and psychological worlds (Tausczik and Pennebaker 2010). Styles encompass a range of linguistic features, including sentence structure, grammar, and punctuation patterns. Unlike the content, which continuously changes and can be influenced by subject matter, external sources, or intentional deception, stylistic features are intrinsic to one’s writing and are less prone to deliberate manipulation (da Rocha et al. 2022).
4.1 Identify Extreme Groups using Styles
Here, we answer RQ2 by proposing a style classifier and demonstrating that we can distinguish ordinary online communities from different online extreme groups based on style alone.
Lee et al. (2022) showed that political groups with different ideologies exhibit distinct tendencies when consuming and disseminating information on social media platforms. In order to investigate styles of ideological online groups and examine if different groups show different text styles, we use the two collections of Facebook groups that represents two datasets of extreme groups – antivax and far-right, as detailed in Section 2.
@@ 4.1 Identify Extreme Groups using Styles
We begin by demonstrating the effectiveness of using style to distinguish between extreme groups and a “benign” control group. We use the scikit-learn Python library [^5] https://scikit-learn.org/stable/modules/multiclass.html for all classifiers used in this section; Logistic regression (LR), Linear SVC (SVC), and Random Forest (RF). All three classifiers inherently support multiclass classification, and we use the default strategy for each classifier.
We design a predictive experiment aimed at evaluating and contrasting the effectiveness of three dictionary-based stylistic metrics (LIWC, Grievance and StyloMetrix). The task is to distinguish the posts among the three Facebook groups; two extreme Facebook groups and a normal (non-extreme) group. We created the normal group to encompass a wide range of discussions, from cooking to non-profit organizations. To put the performance of stylistic classifiers into context, we compare against a content-based baseline – the popular text encoding technique, BERT (Devlin et al. 2018) commonly used in textual classification tasks. In contrast to the three dictionary-based encodings, BERT considers the contextual information for each instance of a given word, enhancing its capabilities. We use spacy-sentence-bert Python library [^6] https://github.com/MartinoMensio/spacy-sentence-bert to produce vectors of given text. The pretrained model of en_stsb_distilbert_base is used which transforms each sentence into a vector of $768$ dimensions. spacy-sentence-bert is a wrapper of sentence-BERT package which uses mean-pooling strategy by default (Reimers and Gurevych 2019). This approach captures contextual information from the entire input sentence and produces fixed-length vectors suitable for downstream tasks.
Table 5: Performance of five feature sets. We distinguish between two extreme groups – Far right and Antivax – and the normal Facebook group. We randomly sample $1,000$ posts for each group. We report accuracy and macro-F1 to compare among the classifiers and features used. We boldfaced the best results using BERT and LGS.
@@ Results.
Since BERT leverages the content and style of the given text, it is often the best-performing feature set. However, some stylistic-based classifiers outperform the content-based classifiers: the Random Forest classifier, LGS, outperforms BERT. In other words, based solely on the style and without leveraging the content of a given post, we can predict which ideological group the post came from. The true positive rate for the normal group was $0.95$, $0.79$ for the Far right group, and $0.66$ for the Antivax group. We notice that some posts from Antivax group were misclassified as Far right, while only a few posts from Far right were misclassified as Antivax. This is because the anti-vax space is a unique niche with far-right elements, while far-right groups generally discuss diverse topics.
@@ 4.2 Identifying Styles
Here, we identify and classify the writing styles of posts used by people in fringe Facebook communities. This predictive exercise differs from the one in Section 4.1; here, we test whether human-labeled styles can be detected using a style-based classifier; we distinguished user groups in Section 4.1.
@@
First, a team member with a writing background manually annotated the writing style of $100$ text samples from the Far right and Antivax groups ($50$ for each group). She employed a deep qualitative approach in which the style labels were generated alongside the labeling. We identified styles that are extensively used in misinformation online environments by manually inspecting them. The labeled styles in these samples are “Casual”, “Empowerment”, “Clickbait”, “Expert” and “Intimacy”. This annotation is a result of sampling 50 posts from both Far right and Antivax then we remove the posts with no style; Table 6 shows the number of posts per style.
@@
Table 7: Classification results of the three styles; “Casual”, “Empowerment” and “Expert”. A Random Forest (RF) classifier and a Logistic Regression (LR) was used with a stratified split. The results show the average performance of 2-fold cross-validation due to the size of the samples.
@@
Empowerment RF 0.60 0.58 0.51 0.54
LR 0.49 0.52 0.67 0.66
Casual RF 0.67 0.5 0.54 0.55
LR 0.45 0.52 0.67 0.70
Expert RF 0.47 0.5 0.42 0.5
LR 0.40 0.73 0.57 0.46
“Clickbait” and “Intimacy” have only two and one exemplars, respectively, and we decided to remove them from the rest of this analysis. As a result, we classify solely the styles “Casual”, “Empowerment” and “Expert”. We performed a binary classification for each style using One-Versus-Rest (OvR) strategy by randomly sampling negative samples to balance the sample size. For the Random Forest classifier, we used max tree depth = $3$ and the number of trees = $8$ due to the small sample size.
@@ Results.
Table 7 reports the classification results. In general, the stylistic features (LGS) perform better with the Random Forest classifier; BERT shows better performance with the Logistic Regression. This is expected since in BERT embeddings, each dimension is not inherently interpretable or separable and thus not ideal for the Random Forest classifier which excels with features that individually carry strong predictive signals. While it is true that the Random Forest classifier may not fully leverage the rich contextual information in BERT embeddings, we aimed to assess whether it could still yield valuable insights or competitive results when compared to more complex models. This experimental choice was made to offer a comparative perspective on different feature representations and classifiers.
For all three styles, the LGS stylistic features outperform BERT with the Random Forest classifier measured by macro F1 score, which is consistent with the results presented in Table 5 for group style classification. “Expert” style was the most challenging for both classifiers since there were only $13$ labeled samples.
4.3 Identify Far-right articles using Styles: production vs. consumption
Here, we answer RQ3 by distinguishing production from sharing. Inspired by the findings in Section 3.3, we hypothesise that far-right production can be differentiated from far-right consumption on the basis of style. We train the style classifier to distinguish far-right production from far-right consumption, and add moderate production for comparison purposes.
@@
We use the $155,669$ articles shared by far-right Twitter users (see Section 2) as far-right consumption data (FRCons) since far-right Twitter users shared these. In order to increase the number of articles produced by far-right publishers in our sample, as well as the number of moderate articles, we expand the number of articles as follows. Each article shared by Twitter users is added to either the moderate or far-right category based on the leaning of the publisher, as long as the article’s publisher is referenced in allsides data. After this expansion, we have $155,669$ Twitter articles shared by far-right users on Twitter, (FRCons), $41,995$ articles from moderate producers (MDProd), and $14,758$ far-right produced articles (FRProd).
Figure 2: Venn diagram showing the intersection of articles produced (FRProd, MDProd) and consumed (FRCons).
Figure 2 shows the intersection size between these three sets of articles as a Venn diagram. $91\%$ of all far-right produced articles are shared by the far-right consumers, whereas only $39\%$ of the moderate articles are shared by the far-right users. Proportionally, the intersection between far-right produced (FRProd) and far-right shared (FRCons) is larger than the intersection between moderate-produced (MDProd) and far-right shared (FRCons). This shows that, while far-right users opportunistically link to articles from both far-right and moderate producers, they prefer far-right producers.
We build a textual classifier that uses the text of each article and predicts whether it is far-right produced, moderate produced or far-right shared. We downsample each class to the smallest class size ($14,758$ articles).
@@
In Section 4.2, we observed that our curated stylistic features (LGS) perform better with the Random Forest classifier while BERT features outperform with Logistic Regression in most cases for identifying styles.
Figure 3: Classification results of the moderate-produced articles (MDProd), far-right produced articles (FRProd) and the articles consumed by far-right users (FRCons). A Random Forest classifier and a Logistic Regression was used to report macro F1-score from 10-fold stratified cross-validation. The error bars represent 95% confidence intervals. We use ‘lbfgs’ solver with ‘multi_class’ option set to ‘multinomial’ to support three class classification for Logistic Regression.
In Figure 3, we see that LGS with the Random Forest classifier consistently outperforms BERT when distinguishing production from consumption. In addition, LGS with Logistic Regression, exempting for articles from moderate publishers, outperforms BERT as well. Particularly, LGS outperforms BERT by far when distinguishing the far-right consumption group (FRCons). This indicates that it is the style, not the content, of the articles that better characterizes the far-right consumption patterns.
Figure 4: Confusion matrix of the three classes from Random Forest classifier using LGS features.
Figure 4 reports the confusion matrix of the three class classification from the Random Forest classifier. This result is the average of 10-fold stratified cross-validation. The most notable observation is that far-right-produced articles (FRProd) are easily distinguished from the others and far-right consumption (FRCons). The most confusions occurred between the far-right consumption (FRCons) and the moderate produced (MDProd). This result indicates that while far-right production (FRProd) and consumption (FRCons) are clearly separable, the styles utilized by moderate production (MDProd) and far-right consumption (FRCons) are similar. Also, since we did not find a significant difference between far-right and moderate articles in terms of their content (Section 3.2), it is implied that the styles are better signals for information consumers.
These findings reinforce the previous observations that far-right production and far-right consumption patterns are different. More specifically, far-right users do not exclusively consume far-right content but rather cherry pick articles that pursue certain styles. Indeed, style could be a contributing factor driving political polarisation in and of itself. Most notably, the findings imply that the distinctiveness of far-right texts compared to moderate sources is their style; and from this, it follows that perhaps a more distinctive approach to styling moderate views, tailored to different audiences, could help moderate perspectives compete against fringe views in today’s attention economy.